OpenStudy (anonymous):

Solve for m: w = m - 5L A) m = w - 5L B) m = w + 5L C) m = (w - 5)/L D)m = w/L + 5

OpenStudy (texaschic101):

try adding 5L to both sides and what do you get ?

OpenStudy (anonymous):

so it would be w-5L?

OpenStudy (anonymous):

I mean +

OpenStudy (texaschic101):

w = m - 5L ....(add 5L to both sides) + 5L + 5L -------------- w + 5L = m correct....m = w + 5L

OpenStudy (anonymous):

Thank you

OpenStudy (texaschic101):

sure thing :)